The function š‘“(š‘„)=2š‘š‘œš‘  š‘„āˆ’3 is defined for the domain 0ā‰¤š‘„ā‰¤šœ‹/2

a. Find the range of š‘“(š‘„)

b. Find š‘“^āˆ’1(š‘„).


1
Expert's answer
2022-03-17T18:39:11-0400

(i)

F(x)=2cos(x-3)

We begin by finding the magnitude of the trigĀ termĀ 

2cos(x-3)

Ā by taking theĀ absolute valueĀ of theĀ coefficient.

=2


The lower bound of theĀ rangeĀ forĀ cosineĀ is found by substituting the negative magnitude of theĀ coefficientĀ into theĀ equation.


y=-2

The upper bound of theĀ rangeĀ forĀ cosineĀ is found by substituting the positive magnitude of theĀ coefficientĀ into theĀ equation.

y=2

TheĀ rangeĀ isĀ 

-2ā‰¤yā‰¤2


IntervalĀ Notation:

[-2,2]

Set-Builder Notation:

{y|āˆ’2ā‰¤yā‰¤2}


Hence the range becomes,

Range:Ā [āˆ’2,2], {y|āˆ’2ā‰¤yā‰¤2}


(ii)

"f^{-1}(x)"

"f(x)=2cos(x-3)"

"y=2cos(x-3)"

Interchanging x and y,

"x=2cos(y-2)"

Solve for y

"\frac{x}{2}=cos(y-2)"

"y-2=cos^{-1}(\frac{x}{2})"

"y=cos^{-1}(\frac{x}{2})+2"